OpenStudy (anonymous):

if sec theta +tan theta=p,express sec theta,tan theta,sin theta in terms of p

OpenStudy (anonymous):

I think we have to use: \[\huge \color{green}{\sec^2 \theta - \tan^2 \theta = 1}\]

OpenStudy (anonymous):

Or, \[(\sec \theta - \tan \theta)(\sec \theta + \tan \theta) = 1\]

OpenStudy (anonymous):

So, \[\sec \theta - \tan \theta = \frac{1}{p}\]

OpenStudy (anonymous):

Simply add them: \[2\sec \theta = p + \frac{1}{p} \implies \frac{p^2+1}{p}\] So, \[\sec \theta = \frac{p^2 + 1}{2p}\]

OpenStudy (anonymous):

Now subtract them you will get tan\(\theta\)...